OpenStudy (anonymous):

Multiply. Simplify if possible (√6-√8)(√6+√8)

OpenStudy (anonymous):

6-8= -2

OpenStudy (anonymous):

\[(a+b) (a-b) = a^2 - b^2\]

OpenStudy (saifoo.khan):

-2

OpenStudy (anonymous):

can anyone break it down for me? Thank you

OpenStudy (anonymous):

It's a difference of squares. a² - b² = (a + b)(a - b).

OpenStudy (anonymous):

Just Foil it if you don't know the property of difference of squares

OpenStudy (saifoo.khan):

As money bird said we will use the formula, \((a+b) (a-b) = a^2 - b^2\). So we take a as \(\sqrt6\) and b as \(\sqrt8\) inserting the values in the equation, \[(\sqrt6)^2 - (\sqrt 8)^2\]\[6 - 8\]\[-2\]

OpenStudy (anonymous):

Thanks! saifoo.khan

OpenStudy (saifoo.khan):

You welcome.

OpenStudy (saifoo.khan):

Please hit good answer. Thanks.